A measuring cup contains 1/5 of a cup of orange juice. It is then filled to the 1 cup mark with a mixture that contains equal amounts of orange, grapefruit, and pineapple juices. What fraction of the final mixture is orange juice?
Answer by stanbon(75887)

A measuring cup contains 1/5 of a cup of orange juice. It is then filled to the 1 cup mark with a mixture that contains equal amounts of orange, grapefruit, and pineapple juices. What fraction of the final mixture is orange juice?
----------
1st cup DATA:
amount = (1/5) cup ; amount of orange juice = (1/5) cup
--------------
2nd cup DATA:
amount = 1 cup ; Amount of orange juice = (1/3) cup
---------------
Final Mixture DATA:
Amount = (6/5) cup ; amount of orange juice = [(1/5) + (1/3)] = 8/15 cup
-------------------
What fraction of the final mixture is orange juice?
Final answer: (8/15)/(6/5) = (8/15)*(5/6) = 4/9 ot the final mixture is OJ.
==================
Cheers,
Stan H.